# Geospatial analysis of toponyms in geotagged social media posts

### v1

@@ Acknowledgements.
T.H. acknowledges the computational resources provided by the Aalto Science-IT project. N.F. was supported by JSPS KAKENHI Grant Number 24K03007 and JST PRESTO Grant Number JPMJPR21RA, Japan. Data availability: All data and code necessary to reproduce the findings of this paper are deposited and publicly available at https://doi.org/10.5281/zenodo.13860969 and https://github.com/takayukihir/geotagged-tweets. Due to the Twitter’s terms of service, we are unable to redistribute the raw text of the posts we have obtained from the Twitter API. Instead, the data on the number of geotagged posts aggregated at the level of basic grid square cells are available in the above repository. The resident and employed population data are based on 2015 Population Census and 2016 Economic Census for Business Activity, respectively, and are available from the Statistics Bureau of Japan (https://www.stat.go.jp/english/data/kokusei/2015/summary.html, https://www.stat.go.jp/english/data/e-census/2016/outline.html).
@@
S1 Definition of relative entropy
@@
Each distribution is defined in the two dimensional space of the total number of posts per cell, denoted by $n_{\mathrm{all}}$, and the number of posts with word $w$, denoted by $n_{w}$. Specifically, we only consider the values of $n_{\mathrm{all}}$ that are empirically observed, and the values of $n_{w}$ equal to or smaller than $n_{\mathrm{all}}$. That is, the support of each distribution is given by
@@
In Fig S1A, we present the estimated relative entropy for each toponym and each noun. As expected, toponyms denoting cities (such as “Nagoya”) and some submunicipal districts (such as “Shinjuku”), as well as place nouns (such as “park”), are characterized by relatively large values of relative entropy than those for common nouns without a place connotations (e.g., “wallet”), implying that the empirical distribution is highly dissimilar from the location-independent binomial model. However, the relative entropies for toponyms that represent regions (e.g., “Shikoku”), prefectures (e.g., “Ishikawa”), and some other submunicipal districts (e.g., “Setagaya”) are as small as those for common nouns. those for common nouns without a place connotations (e.g., “wallet”). This seems inconsistent with what we expect from S3 Fig in Supporting Information, in which the difference between the data and the model is clearly visible for all domestic toponyms. This is presumably because the data contain more grid cells without any toponym occurrence, i.e., $n_{w}=0$, than expected by the model. Such cells are prevalent even in cells with large $n_{\mathrm{all}}$, in which the model predicts a low probability for $n_{w}=0$. The increase in relative entropy due to the presence of such zero-occurrence grid cells will be more pronounced for words with a large $\hat{p}_{w}$, or, equivalently, a large number of posts $N_{w}$. This hypothesis is supported by the strong positive correlation between $N_{w}$ and $D_{\mathrm{KL}}(Q_{w}\parallel P_{w})$, as shown in Fig S1B. Simply put, a large value of relative entropy may just be an artifact of the popularity of the word.
@@
Figure S1: Dissimilarity of empirical data from the location-independent model, evaluated by relative entropy. (A) Relative entropy $D_{\mathrm{KL}}(Q_{w}\parallel P_{w})$ for each toponym $w$ and noun studied in this work. For each word, 20000 grid cells are randomly sampled 50 times to compute the mean and 95% confidence intervals, indicated by error bars. (B, C) Relative entropy $D_{\mathrm{KL}}(Q_{w}\parallel P_{w})$ and the modified version $D_{\mathrm{KL}}(\tilde{Q}_{w}\parallel P_{w})$, plotted against the total number of posts $N_{w}$ containing word $w$.
@@
To discount the effect of inflated zero occurrences, we only sample grid cells with non-zero occurrences for computing the empirical distribution and relative entropy, i.e., $\Gamma\subseteq\{c\mid n_{w,c}>0\}$. Let $\tilde{Q}$ denote the empirical distribution thus obtained. As shown in Fig 5 in the main text, this approach clearly discriminates the domestic toponyms and place nouns from nonspatial nouns and foreign toponyms. The correlation between $N_{w}$ and $D_{\mathrm{KL}}(\tilde{Q}_{w}\parallel P_{w})$ is much weaker (Fig S1C), further corroborating the hypothesis that the value of $D_{\mathrm{KL}}(Q_{w}\parallel P_{w})$ is dominated by the contribution from zero-occurrence grid cells.

### v2

@@ Acknowledgements.
T.H. acknowledges the computational resources provided by the Aalto Science-IT project. N.F. was supported by JSPS KAKENHI Grant Number 24K03007 and JST PRESTO Grant Number JPMJPR21RA, Japan. Data availability: All data and code necessary to reproduce the findings of this paper are deposited and publicly available at https://doi.org/10.5281/zenodo.13860968 and https://github.com/takayukihir/geotagged-tweets. Due to the Twitter’s terms of service, we are unable to redistribute the raw text of the posts we have obtained from the Twitter API. Instead, the data on the number of geotagged posts aggregated at the level of basic grid square cells are available in the above repository. The resident and employed population data are based on 2015 Population Census and 2016 Economic Census for Business Activity, respectively, and are available from the Statistics Bureau of Japan (https://www.stat.go.jp/english/data/kokusei/2015/summary.html, https://www.stat.go.jp/english/data/e-census/2016/outline.html).
@@
Appendix A Definition of relative entropy
@@
Each distribution is defined in the two-dimensional space of the total number of posts per cell, denoted by $n_{\mathrm{all}}$, and the number of posts with word $w$, denoted by $n_{w}$. Specifically, we only consider the values of $n_{\mathrm{all}}$ that are empirically observed, and the values of $n_{w}$ equal to or smaller than $n_{\mathrm{all}}$. That is, the support of each distribution is given by
@@
In Fig SA1A, we present the estimated relative entropy for each toponym and each noun. As expected, toponyms denoting cities (such as “Nagoya”) and some submunicipal districts (such as “Shinjuku”), as well as place nouns (such as “park”), are characterized by relatively larger values of relative entropy than those for common nouns without a place connotations (e.g., “wallet”), implying that the empirical distribution is highly dissimilar from the location-independent binomial model. However, the relative entropies for toponyms that represent regions (e.g., “Shikoku”), prefectures (e.g., “Ishikawa”), and some other submunicipal districts (e.g., “Setagaya”) are as small as those for common nouns. This seems inconsistent with what we expect from Fig S3 in Supporting Information, in which the difference between the data and the model is clearly visible for all domestic toponyms. This is presumably because the data contain more grid cells without any toponym occurrence, i.e., $n_{w}=0$, than expected by the model. Such cells are prevalent even in cells with large $n_{\mathrm{all}}$, in which the model predicts a low probability for $n_{w}=0$. The increase in relative entropy due to the presence of such zero-occurrence grid cells will be more pronounced for words with a large $\hat{p}_{w}$, or, equivalently, a large number of posts $N_{w}$. This hypothesis is supported by the strong positive correlation between $N_{w}$ and $D_{\mathrm{KL}}(Q_{w}\parallel P_{w})$, as shown in Fig SA1B. Simply put, a large value of relative entropy may just be an artifact of the popularity of the word.
@@
Figure SA1: Dissimilarity of empirical data from the location-independent model, evaluated by relative entropy. (A) Relative entropy $D_{\mathrm{KL}}(Q_{w}\parallel P_{w})$ for each toponym $w$ and noun studied in this work. For each word, 20000 grid cells are randomly sampled 50 times to compute the mean and 95% confidence intervals, indicated by error bars. (B, C) Relative entropy $D_{\mathrm{KL}}(Q_{w}\parallel P_{w})$ and the modified version $D_{\mathrm{KL}}(\tilde{Q}_{w}\parallel P_{w})$, plotted against the total number of posts $N_{w}$ containing word $w$.
@@
To discount the effect of inflated zero occurrences, we only sample grid cells with non-zero occurrences for computing the empirical distribution and relative entropy, i.e., $\Gamma\subseteq\{c\mid n_{w,c}>0\}$. Let $\tilde{Q}$ denote the empirical distribution thus obtained. As shown in Fig 5 in the main text, this approach clearly discriminates the domestic toponyms and place nouns from nonspatial nouns and foreign toponyms. The correlation between $N_{w}$ and $D_{\mathrm{KL}}(\tilde{Q}_{w}\parallel P_{w})$ is much weaker (Fig SA1C), further corroborating the hypothesis that the value of $D_{\mathrm{KL}}(Q_{w}\parallel P_{w})$ is dominated by the contribution from zero-occurrence grid cells.
Appendix B Data preprocessing
In this section, we provide the details of the data preprocessing procedures.
Selection of posts based on source applications
As described in the main text, we limit the sample to posts sent from one of the official or general-use third-party mobile applications. Specifically, posts that have ‘source’ metadata as one of the following applications are included: ‘Twitter for iPhone’, ‘Twitter for Android’, ‘Instagram’, ‘Twitter for iPad’, ‘Twitter for Android Tablets’, ‘Path’, ‘Mobile Web’, ‘Mobile Web (M5)’, ‘Path 2.0’, ‘{CJK*}UTF8minついっぷる for iPhone’, ‘Photos on iOS’, ‘Twitter for Windows Phone’, ‘{CJK*}UTF8minついっぷる Pro for iPhone’, ‘Twitter for BlackBerry®’, ‘Camera on iOS’, ‘{CJK*}UTF8minついっぷる for Android’, ‘{CJK*}UTF8minついっぷる Pro for Android’, ‘{CJK*}UTF8minついっぷる for Android org’, and ‘{CJK*}UTF8minついっぷる Pro for Android org’. {CJK*}UTF8minついっぷる (Twipple) was a Twitter client provided by one of the major Japanese internet service providers, BIGLOBE, until it was discontinued in 2017.
Unnatural geotag distribution
During the data cleaning process, we noticed that the geotags attached to some posts had an unnatural geographic distribution, likely due to manipulation. Figure SA2 shows the geographic distribution of posts containing different numbers of mentions (references to other accounts). There are several rectangular blocks clearly visible in the distributions of posts containing mentions of eight other accounts and nine or more accounts. These blocks are unlikely to be the result of organic distributions for two reasons: (1) some of them extend over sea areas, and (2) such patterns are absent in posts with fewer than seven mentions.
Figure SA3 provides further evidence of manipulation. When posts are categorized by their mention count, the number of posts in each category decreases monotonically with the number of mentions. However, the number of unique grid cells where posts in each category are distributed shows an increase at a mention count of seven. Although the geographic distribution of posts with seven mentions does not exhibit visually clear symptoms of manipulation on the map (Fig SA2), we opted to exclude all posts with seven or more mentions from the dataset as a conservative choice.
Toponyms that are substrings of other toponyms
To exclude references to other toponyms from each toponym subsample, we identify the names of regions, prefectures, cities with a population larger than 50 0005000050\,00050 000, and wards that contain each of the 24 toponyms we study in this work as a substring as follows:
  • {CJK*}
    UTF8min関東 (Kanto) is a substring of {CJK*}UTF8min北関東 (Kitakanto) and {CJK*}UTF8min南関東 (Minamikanto).
  • {CJK*}
    UTF8min四国 (Shikoku) is a substring of {CJK*}UTF8min四国中央 (Shikokuchuo).
  • {CJK*}
    UTF8min九州 (Kyushu) is a substring of {CJK*}UTF8min北九州 (Kitakyushu).
  • {CJK*}
    UTF8min宮城 (Miyagi) is a substring of {CJK*}UTF8min宮城野 (Miyagino).
  • {CJK*}
    UTF8min京都 (Kyoto) is a substring of {CJK*}UTF8min東京都 (Tokyo-to, Tokyo Metropolis).
  • {CJK*}
    UTF8min名古屋 (Nagoya) is a substring of {CJK*}UTF8min北名古屋 (Kitanagoya).
  • {CJK*}
    UTF8min福岡 (Fukuoka) is a substring of {CJK*}UTF8min上福岡 (Kamifukuoka).
  • {CJK*}
    UTF8min広島 (Hiroshima) is a substring of {CJK*}UTF8min北広島 (Kitahiroshima) and {CJK*}UTF8min東広島 (Higashihiroshima).
Figure SA2: Spatial distributions of posts containing a specific number of mentions. Grid cells to which a non-zero number of posts are tagged are colored in blue.
Figure SA3: Number of posts containing a specific number of mentions, and the number of grid cells to which they are tagged.